If a 22 inch rope is to be cut into two pieces so that one piece is 3 inches longer than twice the other, how long is each piece?
Answer by mananth(16949)

let one piece be x
the other will be 2x+3
Length of rope = 22 in
the sum of the pieces = 22
..
x+2x+3=22
3x+3=22
3x=19
x=19/3
x=6.33 in one part
..
the other will be 2x+3
=2*19/3 +3
=38/3 +3
=47/3 inches
..
CHECK
47/3 + 19/3 = 66/3 =22 in